OpenStudy (anonymous):

Calculate the average speed and average velocity of a complete round-trip in which the outgoing 200km is covered at 90km/s, followed by a one-hour lunch break, and return 200km is covered at 50km/h.

OpenStudy (kropot72):

The total time taken, t, is given by: \[\large t=\frac{200}{90}+1+\frac{200}{50}=\frac{65}{9}\ hours\] The total distance, d, covered is: \[\large d=200+200=400\ km\] The average speed, s, is given by: \[\large s=\frac{d}{t}\ .........(1)\] Now you just need to plug the values of d and t into (1) to find the average speed.